# Deterministic Legal Retrieval: An Action API for Querying the SAT-Graph RAG

### v1

@@ Abstract
The Structure-Aware Temporal Graph RAG (SAT-Graph RAG) addresses core limitations of standard Retrieval-Augmented Generation in the legal domain by providing a verifiable knowledge graph that models hierarchical structure, temporal evolution, and causal events of legal norms. However, a critical gap remains: how to reliably query this structured knowledge without sacrificing its deterministic properties. This paper introduces the SAT-Graph API, a formal query execution layer centered on canonical actions—atomic, composable, and auditable primitives that isolate probabilistic discovery from deterministic retrieval. These actions enable: (i) high-precision hybrid search; (ii) robust reference resolution; (iii) point-in-time version retrieval; and (iv) auditable causal tracing. We demonstrate how planner-guided agents can decompose complex queries into Directed Acyclic Graphs (DAGs) of these actions. This two-layer architecture transforms retrieval from an opaque black box to a transparent, auditable process, directly addressing Explainable AI (XAI) requirements for high-stakes domains.
Keywords: Legal AI, Knowledge Graphs, Retrieval-Augmented Generation (RAG), Agentic Frameworks, Explainable AI, Trustworthy AI, Temporal Data Modeling, Legal Ontology.
@@ 1 Introduction
The advent of Large Language Models (LLMs) has catalyzed a paradigm shift in information retrieval, with Retrieval-Augmented Generation (RAG) emerging as a dominant architecture for grounding model outputs in factual data [1]. By retrieving relevant text chunks to supplement the model’s parametric knowledge, RAG systems have shown remarkable success in reducing hallucinations. However, the direct application of standard "flat-text" RAG to high-stakes, structured domains such as law reveals fundamental limitations. Legal corpora are not mere collections of text; they are complex, interlinked systems defined by a rigid hierarchical structure, a continuous temporal evolution, and a clear causal chain of legislative actions. Recent studies confirm that traditional RAG frameworks, which rely on surface-level semantic similarity, fail to capture the nuanced relevance required by legal statutes and are insufficient for domain-specific applications without significant adaptation [2, 3].
Standard RAG architectures are ill-equipped to handle this complexity. Even when enhanced with sophisticated semantic segmentation that correctly parses a legal text into its constituent articles and paragraphs, the underlying retrieval mechanism remains fundamentally flawed. It treats these well-defined segments as a simple "flat-text" collection. More critically, it lacks any inherent understanding of versioning. A query about a legal provision is meaningless without a temporal anchor, as the correct answer changes over time. This structural and causal blindness renders such systems fundamentally unreliable for applications demanding legal precision and auditability.
To address these deficiencies, our prior work introduced the Structure-Aware Temporal Graph RAG (SAT-Graph RAG) [4]. This framework moves beyond flat text by modeling legal norms as a verifiable knowledge graph whose ontology explicitly represents the hierarchy, diachronic evolution, and causal events of legislation, as illustrated in Figure 1. While the SAT-Graph provides the necessary verifiable knowledge substrate, a critical gap remains: how can external systems, such as agentic reasoning frameworks or query engines, reliably and flexibly interact with this structured knowledge without sacrificing its deterministic properties? Unstructured access could easily bypass the graph’s guarantees, undermining the very trustworthiness it was designed to provide.
@@
This paper bridges that critical gap. We argue that a trustworthy knowledge base requires a formally defined and auditable interaction protocol. To this end, we introduce a formal query execution layer, which we specify as the SAT-Graph API. The core of this API is a library of canonical actions—atomic, composable, and auditable primitives designed to serve as a secure low-level interface for querying the SAT-Graph. Instead of a monolithic query engine, we provide the essential building blocks that higher-level systems can use to construct complex retrieval plans. These actions provide a comprehensive toolkit for: (i) high-precision hybrid search combining semantic, lexical, and structured metadata filters; (ii) robust resolution of textual references to formal identifiers; (iii) deterministic retrieval of point-in-time versions; and (iv) complete, auditable tracing of causal lineage.
We demonstrate how a planner-guided agent, such as those proposed in advanced Graph RAG frameworks [5], can leverage this API by decomposing a user’s natural language query into an executable Directed Acyclic Graph (DAG) of these actions. This two-layer architecture—with the SAT-Graph as the verifiable knowledge base and our Action API as the secure interaction protocol—establishes a new blueprint for building trustworthy legal AI. It is a direct architectural response to the broader challenge of Explainable AI (XAI), moving the system’s reasoning from an opaque "black box" to a transparent process that can be audited and understood, a critical requirement for high-stakes domains [6]. This allows advanced reasoning engines to operate on legal knowledge while ensuring every step of the retrieval process remains explicit, verifiable, and grounded in the structure of the law.
@@
  • We formally specify a Canonical Action API, a library of primitive and composable actions for interacting with legal knowledge graphs, covering primitives for retrieval, temporal resolution, and causal analysis.
@@
  • We articulate how this architecture decouples the knowledge representation (the graph) from the query logic (the agent), enabling a robust and explainable approach to answering legal questions with LLMs.
The remainder of this paper is structured as follows. Section 2 discusses related work in legal AI and graph-based retrieval. Section 3 details the specification of our Canonical Action API. Section 4 presents use cases demonstrating its application, and Section 5 concludes with a discussion of the implications and future directions.
@@ 2.2 Retrieval-Augmented Generation and its Limitations
While effective for unstructured text, this "flat-text" RAG paradigm falters in the legal domain. Recent benchmarks confirm that the retrieval stage remains a critical bottleneck for legal RAG systems [12]. Our claims of "context blindness" and "temporal naivety" are not hypothetical; they reflect well-documented failure modes. Studies show that a reliance on surface-level similarity fails to capture the "nuanced relevance" of legal statutes [3] and that standard RAG is insufficient without domain-specific adaptations to handle the complexity of legal queries [2]. The inability of these systems to guarantee the retrieval of the correct historical version of a law is a fatal flaw for any serious legal application.
@@ 2.3 Knowledge Graphs for Retrieval (Graph RAG)
Our SAT-Graph RAG [4] aligns with this trend but makes a critical distinction. Unlike generalist approaches that discover a graph "bottom-up" from unstructured text, the SAT-Graph is constructed "top-down," mapping the pre-existing, formal structure of the law onto a deterministic, temporal ontology. The most important entities are not named persons or places, but the structural components of the law itself and the legislative events that alter them. While this provides a robust knowledge base, the question of a standardized interaction protocol remains open.
@@ 2.4 Agentic Frameworks for Reasoning and Retrieval
Recent advances have seen LLMs employed not just as generators, but as reasoning agents capable of planning and tool use [15]. These agents can decompose complex questions and utilize a "library of tools" to gather information. The API we propose is, in effect, this specialized tool library for the legal graph.
A critical insight from recent research is the symbiotic relationship between architectural innovation and evaluation benchmarks. The legal domain has historically suffered from a lack of benchmarks that require deep, multi-hop reasoning over structured, temporal data. However, new, more challenging benchmarks are emerging, such as "Bar Exam QA" and "Housing Statute QA," which are specifically designed to require "reasoning-intensive" queries rather than simple lexical overlap [16]. These new benchmarks expose the deep limitations of traditional RAG and create a clear need for architectures like ours. The SAT-Graph, combined with our Canonical Action API, is precisely the type of verifiable, structure-aware framework needed to successfully address these next-generation evaluation tasks.
@@ 2.5 Positioning Our Contribution
While prior work has focused on either the graph structure itself or the high-level agentic logic, our contribution is the formal specification of the intermediate layer: a Canonical Action API. This is not another monolithic retrieval system, but a low-level protocol that provides the necessary safety guarantees for the legal domain. It enables advanced agents to perform complex reasoning while ensuring that every interaction with the knowledge base is explicit, verifiable, and grounded in the temporal and structural reality of the law.
3 Specification of the Canonical Action API
The central thesis of our work is that trustworthy interaction with a structured knowledge base like the SAT-Graph requires a formal, auditable protocol. We introduce this protocol as the Canonical Action API, a library of atomic and composable actions. This API serves as an abstraction layer, hiding the underlying graph database implementation (e.g., Cypher queries over a Neo4j database). Crucially, it provides a unified interface for both probabilistic discovery actions (like resolving natural language references) and fully deterministic retrieval actions that operate on formal identifiers. In this pattern, an external agentic framework is designed to interact with the knowledge base primarily through these actions, ensuring that every step in a retrieval plan is explicit, verifiable, and its level of certainty is understood, as shown in Figure 2.
Figure 2: Diagram illustrating a reasoning agent decomposing a user prompt into tasks that are executed through actions provided by the SAT-Graph API.
@@ 3.1 Design Principles
The design of the API is guided by three core principles that, together, ensure trustworthiness even when interfacing with probabilistic language models.
  • Maximal Determinism: We acknowledge that any system interpreting natural language has inherently probabilistic components. Our design philosophy is to isolate this non-determinism at the API’s entry points. Actions for discovery and search are probabilistic by nature. However, the core architectural contribution of this API is to enable the orchestration of fully deterministic retrieval plans once a formal identifier is obtained. It is this capability for achieving determinism in the final, grounded query that fundamentally distinguishes our approach and justifies the "Deterministic Retrieval" paradigm. Once a formal id (URI/URN) is acquired, all subsequent actions that operate on it are guaranteed to be deterministic.
  • Composability: Actions are designed as atomic "building blocks." They can be chained together by a planning agent to construct complex query workflows. This allows for flexibility while ensuring that the overall plan is composed of individually verifiable steps, clearly distinguishing between initial probabilistic grounding and subsequent deterministic traversal. Server-side execution of composed actions ensures efficient processing even for complex multi-step plans.
  • Verifiability through Auditability: This principle is our safeguard against the initial non-determinism. Every action, including the probabilistic ones, must return not just a result but also a confidence score or justification. When an agent executes a plan, the sequence of action calls and their outputs forms a complete, human-readable audit trail. This log makes the entire process verifiable, even if not every step is deterministic. A user can inspect the log and see, "The system resolved ’Article 6’ to id ’…’ with 98% confidence," allowing them to trust or question the initial grounding step.
@@ 3.2 Core Data Models
The actions in our API operate on and return a set of core data objects that correspond directly to the entities in the SAT-Graph ontology, as illustrated in Figure 3. The ontology makes a clear distinction between versioned normative entities (Item) and atemporal conceptual entities (Theme).
@@
  • id (ID): A unique, canonical identifier (e.g. URI/URN).
  • category (string): The architectural class (‘Work‘ or ‘Work Component‘).
  • type (string): The specific semantic subtype (‘Constitution‘, ‘Article‘, ‘Chapter‘).
@@
  • metadata (JSON)?: (Optional) A flexible json store for structured, atemporal properties of the item—examples of metadata description of a legal norm’s Work in [17].
@@
Represents an atemporal, conceptual entity used to classify Items and to organize knowledge. Themes can be arranged in their own taxonomy.
@@
  • parent (list[id])?: (Optional) A list of ids of broader themes. Themes can have multiple parents, creating a directed acyclic graph (DAG) structure. Root themes have an empty list.
@@
A foundational design principle of this framework is the strict separation of normative entities (Item) from conceptual, classificatory entities (Theme). Items represent the formal, versioned structure of the law itself. Themes, in contrast, form an independent, atemporal ontology used to organize and discover knowledge. An Item’s structural relationships (‘parent‘, ‘children‘) define its fixed position within a document, forming a strict hierarchy. A Theme’s relationships define both its taxonomic position relative to other themes (‘parent‘, ‘children‘) and its classificatory relationship to the law (‘members‘). This clear separation allows an agent to navigate the formal structure of a norm (or component) and the conceptual map of the legal domain through distinct, unambiguous API actions.
@@
Represents a specific, time-bound version of an Item. An Version captures not only the content but also the structural position of an item at a specific point in time.
@@
  • item_id (ID): The id of the abstract Item it is a version of.
@@
  • parent (list[ID])?: (Optional) A list of identifiers for its parent Versions. A ‘Work‘’s Version has no parents.
  • children (list[ID])?: (Optional) A list of identifiers for its children Versions, representing the complete structure at this point in time.
@@
Represents a reified legislative event as the most granular, atomic unit of change in the graph. It is a first-class entity that models a single, auditable causal link, connecting a specific clause of an authorizing law to the specific change it enacts.
@@
  • source_version_id (ID): The identifier for the specific Version of the legal text that authorizes this single change.
@@
  • produces_version (ID): The specific Version that is created by this action.
@@
Represents a concrete piece of textual information. It is a flexible object used to hold any "aspect" of text associated with any primary entity in the graph. This could be the canonical text of a versioned item (an Version), a detailed description of a Theme, or a textualized metadata aspect of an Item (like a publication info.).
@@
  • source_node_type (string): The type of the source node (e.g., "Item", "Version", "Action", "Theme"), to allow for unambiguous retrieval by an agent.
@@
3.3 API Action Specification
The API is organized into functional categories based on agentic intent. A core design guarantee governs all actions: any action that accepts a formal ID as its primary input is fully deterministic. The only exceptions are the discovery and search actions, which interpret natural language text. This section details the signature and purpose of each action.

### v2

# Deterministic Legal Agents: A Canonical Primitive API for Auditable Reasoning over Temporal Knowledge Graphs

@@ Abstract
For autonomous legal agents to operate safely in high-stakes domains, they require a foundation of absolute determinism and auditability—guarantees that standard Retrieval-Augmented Generation (RAG) frameworks cannot provide. When interacting with temporal knowledge graphs that model the complex evolution of legal norms, agents must navigate versioning, causality, and hierarchical structures with precision, a task for which black-box vector search is ill-suited. This paper introduces a new architectural pattern to solve this: a formal Primitive API designed as a secure execution layer for reasoning over such graphs. Instead of a monolithic query engine, our framework provides a library of canonical primitives—atomic, composable, and auditable primitives. This design empowers planner-guided agents to decompose complex legal questions into transparent execution plans, enabling critical tasks with full verifiability, including: (i) precise point-in-time version retrieval, (ii) robust causal lineage tracing, and (iii) context-aware hybrid search. Ultimately, this architecture transforms opaque retrieval into auditable reasoning, turning the agent’s internal process from a black box into a verifiable log of deterministic primitives and providing a blueprint for building the next generation of trustworthy legal AI.
Keywords: Legal AI, Deterministic Legal Agents, Agentic Frameworks, Temporal Knowledge Graphs, SAT-Graph, Auditability, Explainable AI (XAI), Trustworthy AI, Retrieval-Augmented Generation (RAG).
@@ 1 Introduction
The emergence of agentic frameworks powered by Large Language Models (LLMs) has catalyzed a new paradigm in automated reasoning, with Retrieval-Augmented Generation (RAG) serving as a dominant architecture for grounding agent outputs in factual data [1]. While RAG helps reduce hallucinations, the direct application of its standard "flat-text" model by agents in high-stakes, structured domains such as law reveals fundamental limitations. Legal corpora are not mere collections of text; they are complex, interlinked systems defined by a rigid hierarchical structure, a continuous temporal evolution, and a clear causal chain of legislative actions. Recent studies confirm that traditional RAG frameworks, which rely on surface-level semantic similarity, are insufficient for building reliable, domain-specific agents without significant adaptation [2, 3].
Standard RAG architectures are ill-equipped to provide the guarantees these agents require. Even when enhanced with sophisticated semantic segmentation, the underlying retrieval mechanism remains fundamentally flawed. It treats well-defined legal segments as a simple "flat-text" collection. More critically, it lacks any inherent understanding of versioning. A query about a legal provision is meaningless without a temporal anchor, as the correct answer changes over time. This structural and causal blindness is not merely a technical limitation; it is a critical vulnerability that renders any agent relying on it fundamentally unreliable for applications demanding legal precision and auditability.
To address these deficiencies, our prior work introduced the Structure-Aware Temporal Graph RAG (SAT-Graph RAG) [4]. This framework moves beyond flat text by modeling legal norms as a verifiable knowledge graph, as illustrated in Figure 1. While the SAT-Graph provides the necessary verifiable knowledge substrate, a critical gap remains: how can an autonomous agent reliably and flexibly interact with this structured knowledge without sacrificing its deterministic properties? Unstructured access could allow the agent to bypass the graph’s guarantees, undermining the very trustworthiness it was designed to provide.
@@
This paper bridges that critical gap. We argue that enabling trustworthy legal agents requires more than a better knowledge base; it demands a formally defined and auditable interaction protocol. To this end, we introduce a formal query execution layer, which we specify as the SAT-Graph API—a canonical primitive interface for querying structure-aware, temporal graphs with deterministic retrieval guarantees. The core of this API is a library of canonical primitives—atomic, composable, and auditable primitives designed to serve as the secure toolset for agentic reasoning. Instead of a monolithic query engine, we provide the essential building blocks that agents can use to construct complex, verifiable retrieval plans. These primitives provide a comprehensive toolkit for: (i) high-precision hybrid search combining semantic, lexical, and structured metadata filters; (ii) robust resolution of textual references to formal identifiers; (iii) deterministic retrieval of point-in-time versions; and (iv) complete, auditable tracing of causal lineage [^1] It is important to clarify the scope of ”causality” within this framework. We address the chain of legislative events that determine the textual validity of a legal provision at a specific point in time (causality in the textual-documentary world). This is distinct from the legal-dogmatic concept of causality, which typically refers to the relationship between a real-world event (the operative fact) and the legal norm that applies to it..
We demonstrate how a planner-guided agent, such as those proposed in advanced Graph RAG frameworks [5], can leverage this API by decomposing a user’s natural language query into an executable Directed Acyclic Graph (DAG) of these primitives. This two-layer architecture—utilizing a temporal knowledge graph like the SAT-Graph as the verifiable knowledge base, and our API as the secure interaction protocol—establishes a new blueprint for building deterministic legal agents. It is a direct architectural response to the broader challenge of Explainable AI (XAI), moving the agent’s reasoning from an opaque "black box" to a transparent process that can be audited and understood, a critical requirement for high-stakes domains [6]. This allows advanced agents to operate on legal knowledge while ensuring every step of the retrieval process remains explicit, verifiable, and grounded in the structure of the law.
@@
  • We formally specify a Canonical Primitive API, a library of deterministic primitives that serve as the necessary toolset for building auditable agents.
@@
  • We introduce a conceptual shift for RAG, transforming the retrieval process from a passive, single-step context injection into an active, agent-driven inquiry where the agent strategically plans and executes the construction of its own context.
  • We articulate how this architecture decouples the knowledge representation (the graph) from the reasoning logic (the agent), enabling a robust and explainable approach to answering legal questions with LLMs.
The remainder of this paper is structured as follows. Section 2 discusses related work in legal AI and graph-based retrieval. Section 3 details the specification of our API. Section 4 presents use cases demonstrating its application, and Section 5 concludes with a discussion of the implications and future directions.
@@ 2.2 Retrieval-Augmented Generation and its Limitations
While effective for unstructured text, this "flat-text" RAG paradigm falters in the legal domain. Recent benchmarks confirm that the retrieval stage remains a critical bottleneck for legal RAG systems [12]. Our claims of "context blindness" and "temporal naivety" are not hypothetical; they reflect well-documented failure modes. Studies show that a reliance on surface-level similarity fails to capture the "nuanced relevance" of legal statutes [3] and that standard RAG is insufficient without domain-specific adaptations to handle the complexity of legal queries [2]. The inability of these systems to guarantee the retrieval of the correct historical version of a legal norm is a fatal flaw that renders any agent built upon them fundamentally untrustworthy and unsafe for any serious legal application.
@@ 2.3 Knowledge Graphs for Retrieval (Graph RAG)
Our SAT-Graph RAG [4] aligns with this trend but makes a critical distinction. Unlike generalist approaches that discover a graph "bottom-up" from unstructured text, the SAT-Graph is constructed "top-down," mapping the pre-existing, formal structure of the law onto a deterministic, temporal ontology. The most important entities are not named persons or places, but the structural components of the law itself and the legislative events that alter them. While this provides a robust knowledge base, the critical question of how an agent can interact with this knowledge deterministically and auditably remains unaddressed. Without a formal protocol, the guarantees of the graph are easily bypassed, reintroducing the very unreliability we seek to eliminate.
@@ 2.4 Agentic Frameworks for Reasoning and Retrieval
Recent advances have seen LLMs employed not just as generators, but as reasoning agents capable of planning and tool use [15]. However, these generic agentic architectures are only as reliable as the tools they are given. In a high-stakes domain like law, a non-deterministic or opaque retrieval tool is a critical point of failure for the entire agentic system. The API we propose is designed to be this missing high-fidelity tool library, providing the deterministic guarantees required for auditable reasoning.
A critical insight from recent research is the symbiotic relationship between architectural innovation and evaluation benchmarks. The legal domain has historically suffered from a lack of benchmarks that require deep, multi-hop reasoning over structured, temporal data. However, new, more challenging benchmarks are emerging, such as "Bar Exam QA" and "Housing Statute QA," which are specifically designed to require "reasoning-intensive" queries rather than simple lexical overlap [16]. These new benchmarks expose the deep limitations of traditional RAG and create a clear need for architectures like ours. The SAT-Graph, combined with our Canonical Primitive API, is precisely the type of verifiable, structure-aware framework needed to successfully address these next-generation evaluation tasks.
@@ 2.5 Positioning Our Contribution
While prior work has focused on either the graph structure itself or the high-level agentic logic, our contribution is the formal specification of the intermediate layer: a Canonical Primitive API. This is not another monolithic retrieval system, but a low-level protocol that provides the necessary safety guarantees for the legal domain. It enables advanced agents to perform complex reasoning by transforming retrieval from a single, probabilistic primitive into a deterministic, auditable plan. Every interaction with the knowledge base becomes explicit, verifiable, and grounded in the temporal and structural reality of the law, providing a true foundation for trustworthy legal agents.
3 Specification of the Canonical Primitive API
The central thesis of our work is that for legal agents to be trustworthy, their interaction with a structured knowledge base must be governed by a formal, auditable protocol. We introduce this protocol as the SAT-Graph API, a library of atomic and composable primitives that serve as the agent’s verifiable toolset. This API serves as an abstraction layer, hiding the underlying graph database implementation (e.g., Cypher queries over a Neo4j database). Crucially, it provides a unified interface for both probabilistic discovery primitives (like resolving natural language references) and fully deterministic retrieval primitives that operate on formal identifiers. In this pattern, an autonomous agent is empowered to interact with the knowledge base exclusively through these primitives, ensuring that every step in its reasoning plan is explicit, verifiable, and its level of certainty is understood, as shown in Figure 2.
Figure 2: Diagram illustrating a reasoning agent decomposing a user prompt into tasks that are executed through primitives provided by the SAT-Graph API.
@@ 3.1 Design Principles
The design of the API is guided by three core principles that, together, ensure the agent’s trustworthiness, even when its core logic is a probabilistic language model.
  • Maximal Determinism: We acknowledge that any system interpreting natural language has inherently probabilistic components. Our design philosophy is to isolate this non-determinism at the API’s entry points. Primitives for discovery and search are probabilistic by nature. However, the core architectural contribution of this API is to enable an agent to orchestrate fully deterministic reasoning plans once a formal identifier is obtained. This capability for achieving determinism is what distinguishes our approach and justifies the "Deterministic Legal Agent" paradigm. Once a formal id (URI/URN) is acquired, all subsequent primitives that operate on it are guaranteed to be deterministic.
  • Composability: Primitives are designed as atomic "building blocks." They can be chained together by a planning agent to construct a verifiable execution plan. The architectural philosophy is to provide the agent with powerful, deterministic primitives, while the reasoning logic that composes them (e.g., iterating through a list, inspecting text) remains the responsibility of the agent. This deliberate separation ensures that the agent’s entire reasoning process—not just the API calls—is explicit and auditable, preventing opaque server-side operations from obscuring the plan.
  • Verifiability through Auditability: This principle is our safeguard against the initial non-determinism. Every primitive, including the probabilistic ones, must return not just a result but also a confidence score or justification. When an agent executes a plan, the sequence of primitive calls and their outputs forms a complete, human-readable audit trail. This log is the tangible artifact of the agent’s reasoning process, making it fully verifiable, even if the initial discovery steps are probabilistic. A user can inspect the log and see, "The system resolved ’Article 6’ to id ’…’ with 98% confidence," allowing them to trust or question the initial grounding step.
@@ 3.2 Core Data Models
The primitives in our API provide an agent with a standardized vocabulary of data objects for perceiving and navigating the knowledge graph. These core objects, which correspond directly to the entities in the SAT-Graph ontology, as illustrated in Figure 3.
@@
  • id (ID): A unique, canonical identifier (e.g. URN).
  • type (id): The specific item type id.
@@
  • metadata (JSON)?: (Optional) A flexible json store for structured properties of the item—examples of metadata description of a legal norm’s Work in [17].
@@
Represents a conceptual entity used to classify Items and to organize knowledge. Themes can be arranged in their own taxonomy.
@@
  • uri (string)?: (Optional) A uniform resource identifier.
  • parent (list[ID])?: (Optional) A list of ids of broader themes. Themes can have multiple parents. Root themes have an empty list.
@@
  • metadata (JSON)?: (Optional) A flexible json store for structured properties of the theme.
@@
A foundational design principle of this framework is the strict separation of normative entities (Item) from conceptual, classificatory entities (Theme). Items represent the formal, versioned structure of the law itself. Themes, in contrast, form an independent, atemporal ontology used to organize and discover knowledge. An Item’s structural relationships (‘parent‘, ‘children‘) define its fixed position within a document, forming a strict hierarchy. A Theme’s relationships define both its taxonomic position relative to other themes (‘parent‘, ‘children‘) and its classificatory relationship to the law (‘members‘). This clear separation is a critical design choice that empowers the agent. It allows the agent to disambiguate its intent, navigating the formal structure of a norm (its hierarchy) and the conceptual map of the legal domain (its themes) through distinct, unambiguous API primitives.
@@
While the underlying ontology distinguishes between different semantic types of hierarchy, the API deliberately abstracts these into a conceptually unified pair of properties: parent and children. This pragmatic choice simplifies an agent’s logic, allowing it to reason about "moving up" or "moving down" a hierarchy using the same vocabulary, regardless of the entity type. We acknowledge a minor difference in the data types: an Item has a single parent, while Themes and Versions can have a list of parents. This is an intentional design choice reflecting the semantics of the graph. An Item’s position in a document’s structure is fixed and singular. In contrast, a Theme can belong to multiple broader categories, and a Version can be reused under multiple parent versions as a performance optimization. Despite this type difference, the conceptual abstraction remains consistent: parent always represents a move towards a broader or containing entity, and children a move towards a more specific or contained one. This approach prioritizes a consistent agentic vocabulary without sacrificing the semantic richness of the underlying model.
@@
Represents a specific, time-bound version of an Item. A Version object is the agent’s key to the past, capturing not only the content but also the structural position of an item at a specific point in time.
@@
  • item (ID): The id of the abstract Item it is a version of.
@@
  • uri (string)?: (Optional) A uniform resource identifier.
  • parent (list[ID])?: (Optional) A list of identifiers for its parent Versions. While a component has one structural parent, this is modeled as a list to support advanced graph optimizations, such as reusing an unchanged child Version under multiple, newly-created parent Versions. A Work’s Version has no parents.
@@
Represents a reified legislative event as the most granular, atomic unit of change in the graph. It is a fundamental building block of the agent’s audit trail, a first-class entity that models a single, verifiable causal link, connecting a specific clause of an authorizing law to the specific change it enacts.
@@
  • source_version (ID): The identifier for the specific Version of the legal text that authorizes this single change.
@@
  • produces_version? (ID): (Optional) The specific Version that is created by this action.
  • metadata (JSON)?: (Optional) A flexible json store for structured properties of the action.
@@
Represents a concrete piece of textual information. It is a flexible object used to hold any "aspect" of text associated with any primary entity in the graph. This could be the canonical text of an Version, a detailed description of a Theme, or a textualized metadata aspect of an Item (like a publication info.).
@@
  • source_node_type (string): The type of the source node (e.g., "Theme", "Item", "Version", "Action"), to allow for unambiguous retrieval by an agent.
@@
3.3 Canonical Primitive API Specification
The SAT-Graph API provides an agent with a structured vocabulary of primitives, organized by intent. A core design guarantee governs this vocabulary: any primitive that accepts a formal ID as its primary input is fully deterministic. This clear distinction is critical: it allows the agent to consciously manage uncertainty during its discovery phase, before transitioning to a fully deterministic and auditable execution path. This section details the signature and purpose of each primitive, which serve as the building blocks for the agent’s reasoning plans.